5) Given: Ba(NO3)2(aq) + Na2SO4(aq) → BaSO4(s) + 2NaNO3(aq) Step 1 is done!
How many grams of NaNO3 are produced from 3.5 g of Ba(NO3)2 ?
Step 2: Mass to moles
Step 3: Mole to Mole (Ratio)
Step 4: Mole to Mass
1
Expert's answer
2020-12-01T14:22:30-0500

Ba(NO3)2(aq) + Na2SO4(aq) → BaSO4(s) + 2NaNO3(aq)

n = m/M

M (Ba(NO3)2) = 261.3 g/mol

M (NaNO3) = 85 g/mol

n (Ba(NO3)2) = 3.5/261.3 = 0.013 mol

n (NaNO3) = 2 x n (Ba(NO3)2) = 2 x 0.013 = 0.03 mol

m (NaNO3) = n x M = 0.03 x 85 = 2.3 g